# CLASH: Collaborative Large-Small Hierarchical Framework for Continuous Vision-and-Language Navigation

### v1

@@ IV-A2 Framework of the RSMP
a. Language Encoding Module: To capture the semantic structure of the instruction $I$, a pretrained language model such as BERT [devlin2019bert] is utilized. Let $\psi(\cdot)$ be the linear transformation layer, $\phi(\cdot)$ the embedding layer, and $\mathcal{F}_{I}(\cdot)$ the transformer encoder. The resulting language feature representation $Q_{s}\in\mathbb{R}^{L\times d_{h}}$ is computed as $Q_{s}=\mathcal{F}_{I}(\psi_{i}(I)+\phi_{i}(P_{I}))$, where $L$ denotes the number of the instruction tokens, $d_{h}$ is the hidden dimension of the feature embeddings, and $P_{I}$ denotes the absolute positional indices of the tokens. When the causal learning module is applied, the text content causal representations become $G_{s}=LN[\psi_{s}(BA(Q_{s},Z_{k}))]$, where $LN$ denotes the layernorm function, and $Z_{k}$ is the textual confounder features.
b. Panoramic Visual Encoding Module: To comprehensively perceive its surroundings, the agent processes a panoramic observation $E$ composed of 12 perspective sub-images. Each sub-image is first encoded by the pre-trained visual encoder $\mathcal{F}_{V}(\cdot)$ (like CLIP [radford2021learning]). The relative orientation of each sub-image is represented by a directional embedding $\gamma_{im}={(\sin\theta_{i},\cos\theta_{i},\sin\beta_{i},\cos\beta_{i})}_{i=1}^{12}$. A multi-layer self-attention encoder ($\mathcal{F}_{\text{SA}}$) then fuses the multi-view features to enable cross-view interaction, obtaining $Q_{v}\in\mathbb{R}^{12\times d_{h}}$:
@@
c. Global-Local Cross-Modal Localization Module: In widely used VLN-CE hierarchical frameworks [an2024etpnav, wang2025g3d, wang2024lookahead], cross-modal fusion is typically performed only at the global level between the instruction and the historical topology representation. However, this design may overlook fine-grained local context. Motivated by the effectiveness of dual local–global architectures in discrete settings [chen2022think, wang2023scaling], we introduce a local cross-modal fusion branch to enhance detailed environmental grounding and action-level decision-making.
@@ IV-B1 Panoramic Visual Prompt (PVP)
To enable holistic spatial understanding, different from prior methods that rely on discrete sub-views or sequential video streams [chen2024mapgpt, zhou2024navgpt2, cheng2024navila], we directly feed the full panoramic observation into the MLLM. Candidate waypoints are projected onto the panorama to visually encode their spatial positions, thereby enhancing the alignment between language instructions and navigational intent. Additionally, we observe that the lateral edges of the equirectangular panorama correspond to the agent’s backward view, which can mislead MLLMs into interpreting rear-view regions as lateral surroundings. Since navigation reasoning primarily depends on the forward field of view, we introduce a semi-transparent mask that attenuates the visual saliency of rear-view regions while preserving global scene continuity. This design reduces distractive cues from the backward view and helps the model concentrate on task-relevant front-view information. Our design is theoretically motivated by recent advances in attention prompting [yu2024attention], which demonstrate that explicit visual prompts or region-wise masking can steer the attention distribution of large vision-language models and improve spatial reasoning. An illustration of PVP and the masking effect is shown in Fig. 3.
@@ IV-B3 Causal chain-of-thought
To improve generalization, we adopt causal learning [wang2024causal] for the RSMP to mitigate dataset bias. Large models, trained on diverse corpora, naturally capture causal relations and require no structural change. Instead, we apply chain-of-thought (CoT) prompting [wei2022chain] to enhance reasoning and interpretability. RLMR is guided to reason through five steps: trajectory status, action planning, visual grounding, suggestion evaluation, and final decision, with an assigned confidence score $\mathbf{p}^{\text{RLMR}}$. The complete prompt template is provided in the appendix.
@@ IV-C Uncertainty-Aware Collaboration Mechanism (UCM)
Existing works demonstrated that RSMP, operating within a panoramic-view hierarchical structure, consistently outperforms RLMR, even when the latter is fine-tuned on the in-domain VLN dataset [wang2024magic, zhou2024navgpt, zhou2024navgpt2]. While RLMR has the great potential to offer greater transparency and reflective reasoning, it is more prone to errors in long-horizon embodied reasoning tasks. To effectively integrate the strengths of both agents, we propose an Uncertainty-Aware Collaboration Mechanism (UCM), which leverages RSMP’s uncertainty estimates to adaptively mediate decision fusion with RLMR.
@@ IV-C1 Conformal Prediction for Uncertainty Quantification
To quantify RSMP’s prediction uncertainty, we adopt conformal prediction (CP) [shafer2008tutorial, huang2024conformal], a distribution-free framework that provides statistically calibrated confidence measures with formal coverage guarantees. CP constructs a prediction set $\mathcal{C}(\mathcal{X})\subseteq\mathcal{Y}$ such that, in expectation over the data distribution, the true label $\hat{y}$ lies within $\mathcal{C}(\mathcal{X})$ with probability at least $1-\epsilon$. This coverage guarantee is statistically derived from the calibration set and is not enforced during evaluation or test inference. Specifically, $\mathcal{X}$ denotes the input observation space (comprising visual observations, language instructions, and navigation history), $\mathcal{C}(\cdot)$ denotes a set of possible output candidates, $\mathcal{Y}$ represents the action label space (possible waypoint candidates), and $\epsilon\in[0,1]$ is a user-specified miscoverage rate (i.e., the allowed probability of the prediction set not containing the true action).
@@ V-A1 Panoramic-RGBD end-to-end waypoint predictor
Most prior hierarchical VLN methods [krantz2022sim, an2024etpnav] adopt an end-to-end waypoint predictor. The model takes RGB $\{v_{i}\}_{i=1}^{12}$ and depth $\{d_{i}\}_{i=1}^{12}$ panoramas, and encodes them using pre-trained visual backbones (denoted as $VE(\cdot)$ and $DE(\cdot)$, respectively). The visual embeddings are then processed by a Transformer encoder ($\mathcal{F}_{\text{SA}}$). A classifier ($\phi_{c}(\cdot)$) outputs a heatmap over a discretized 2D space defined by 120 angles (at $3\degree$ resolution) and 12 distances (from 0.25m to 3.0m in 0.25m increments). Finally, non-maximum suppression (NMS) is applied to the heatmap to extract the top-P candidate waypoints.
@@ V-B1 RGB-D-based point-Goal navigation network
For the local navigation controller, ETPNav [an2024etpnav] uses a “rotate-then-forward” strategy with a Tryout heuristic: if repeated FORWARD actions fail to change position, the agent rotates through preset angles $\Delta\Theta^{try}$ and retries. While effective in simulation, this trial-and-error approach is unsafe for real-world use. It often fails in scenarios with elevation changes, such as approaching a target atop stairs from the side. To address this, we adopt DDPPO [wijmansdd2019ddppo], an end-to-end policy that maps RGB-D inputs and goals to actions, improving obstacle avoidance and deadlock handling in simulation.
@@ V-B2 SLAM-based point-goal navigation algorithm
LiDAR-based SLAM systems are widely used for robust real-world navigation, offering mapping, localization, and path planning in unknown environments without pre-built maps. Although DDPPO performs well in simulation, prior work [gervet2023navigating] shows limited transferability to the real world. To ensure reliable physical deployment, we adopt an online SLAM-based pipeline using ROS move_base for waypoint navigation. This hybrid approach bridges the sim-to-real gap and ensures safe, efficient traversal.
@@ VI-A1 Datasets
The proposed method is evaluated on two widely used VLN-CE benchmarks: R2R-CE and REVERIE-CE, which are adapted from the discrete-path datasets [anderson2018vision, qi2020reverie]. The R2R-CE dataset consists of real-world scanned environments across 90 buildings, with an average length of 32 words. It is split into train (10,819 pairs, 61 scenes), val seen (same buildings as training, 778 pairs and 61 scenes), val unseen (different buildings, 1,839 pairs and 11 scenes), and test unseen (3,408 pairs and 18 scenes). The test set is evaluated via an online server for fair benchmarking. The REVERIE-CE dataset includes 10,318 panoramas from 86 buildings and 21,702 crowd-sourced instructions, averaging 18 words per instruction.
TABLE I: Comparison with other SoTA methods on the R2R-CE dataset [krantz_beyond_2020]. ‘-’ indicates the unavailable statistics. ‘Pano’ denotes whether panoramic visual input is used. $\dagger$ means the usage of the additional data beyond the VLN.
@@
I CM2 [georgakis2022cross] 12.05 6.10 51 43 35 11.54 7.02 42 34 28 13.90 7.70 39 31 24
CWP-CMA [hong2022bridging] 11.47 5.20 61 51 45 10.90 6.20 52 41 36 11.85 6.30 49 38 33
CWP-RecBERT [hong2022bridging] 12.50 5.02 59 50 44 12.23 5.74 53 44 39 13.51 5.89 51 42 36
GridMM [wang2023gridmm] 12.69 4.21 69 59 51 13.36 5.40 57 50 46 13.31 5.55 57 49 45
DreamWalker [wang2023dreamwalker] 11.60 4.09 66 59 48 11.30 5.53 59 49 44 11.80 5.48 57 49 44
ScaleVLN [wang2023scaling] - - - - - - 4.80 - 55 51 - 5.11 - 55 50
ETPNav [an2024etpnav] 11.78 3.95 72 66 59 11.99 4.71 65 57 49 12.87 5.12 63 55 48
BEVBert [an2022bevbert] - - - - - - 4.57 67 59 50 - 4.70 67 59 50
HNR [wang2024lookahead] 11.79 3.67 76 69 61 12.64 4.42 67 61 51 13.03 4.81 67 58 50
UnitedVLN [dai2024unitedvln] - 3.30 78 70 62 - 4.29 70 62 51 - 4.69 68 59 49
g3D-LF [wang2025g3d] - - - - - - 4.53 68 61 52 - 4.78 68 58 51
II InstructNav [long2024instructnav] - - - - - 6.74 6.04 42 30 22 - - - - -
NaVid [zhang2024navid] 10.27 5.47 52 43 38 7.63 5.47 49 37 35 - - - - -
Uni-NaVid [zhang2024uni] 10.04 4.45 65 58 53 9.71 5.58 53 47 43 - - - - -
NaVILA [cheng2024navila] - - - - - - 5.37 58 50 46 - - - - -
NaVILA [cheng2024navila] - 4.82 66 58 53 - 5.22 63 54 49 - - - - -
StreamVLN [wei2025streamvln] - - - - - - 5.43 63 53 47 - - - - -
StreamVLN [wei2025streamvln] - 4.53 68 62 58 - 4.98 64 57 52 - - - - -
@@ VI-A3 Implementation details
Our VLN-CE implementation builds on ETPNav [an2024etpnav]. The RSMP model is initialized with ScaleVLN [wang2023scaling] fine-tuned weights, with CLIP-H/14 [radford2021learning] for image features and learnable backdoor and frontdoor causal modules from GOAT [wang2024causal]. Depth information is incorporated via a pre-trained ResNet-50 [wijmansdd2019ddppo]. Training in Habitat uses a batch size of 16 for 15K iterations with AdamW ($2\times 10^{-5}$), only on the standard R2R training split, and the best checkpoint is selected based on SR and SPL on validation-unseen. Compared with other LLM-based fine-tuning strategies that require large multi-GPU clusters, our method trains efficiently on only one single NVIDIA L40 GPU. REVERIE-CE is fine-tuned from the best R2R-CE checkpoint with identical settings. The RLMR deploys the open-source Qwen2.5-VL-72B [Qwen2.5-VL] locally via VLLM [kwon2023efficient] on 4 A800 GPUs.
@@ VI-B1 Results on R2R-CE
Table I compares CLASH with current SoTA methods on the R2R-CE dataset. For clarity, we categorize existing approaches into three types: Type-I includes conventional end-to-end task-specific models that are fully fine-tuned on the R2R-CE dataset. Type-II represents recent efforts to incorporate large models into VLN-CE. While these large-model approaches perform well in monocula settings, they still fall short of panoramic task-specific models, a trend also observed in discrete VLN benchmarks [zhou2024navgpt, zhou2024navgpt2]. Type-III refers to our proposed collaborative small-large framework, CLASH. It shows that CLASH outperforms existing methods across splits and metrics. Particularly, on the test unseen split, CLASH achieves the 1-st rank in the VLN-CE test leaderboard [^1] https://eval.ai/web/challenges/challenge-page/719/overview, surpassing g3D-LF [wang2025g3d] by a relative 13.79% in SR, and 13.73% in SPL, respectively. Compared to approaches using large vision-language models, CLASH demonstrates substantial gains in both seen and unseen environments. For instance, compared with NaVILA [cheng2024navila], CLASH increases relative SR by 25.86% on val-seen, and 20.37% on val-unseen, respectively.
@@ VI-B2 Results on REVERIE-CE
Table II compares our proposed CLASH method with other SoTA methods on the REVERIE-CE dataset. We mapped the REVERIE coordinates from the discrete environment into the Habitat simulator, enabling data collection, training, and model evaluation. It shows that CLASH outperforms previous methods on all metrics on both val-seen and val-unseen splits. For instance, it shows that on the val-unseen set, CLASH relatively outperforms g3d-LF [wang2025g3d] by 2.92% in SR and 11.76% in SPL, respectively. On the val-seen set, CLASH improves ETPNav [an2024etpnav] by relative 24.94% SR and 24.93% SPL, respectively. This has significantly demonstrated the effectiveness and robustness of our proposed method.
TABLE II: Comparison with other SoTA methods on the REVERIE-CE [qi2020reverie].
@@
Seq2Seq [krantz_beyond_2020] 9.96 22.8 13.7 10.9 10.10 18.6 9.6 7.9
CMA-TF [li2022reve-ce] 9.83 23.2 14.2 11.9 10.27 16.8 8.8 7.3
CMA-SF [li2022reve-ce] 9.84 30.4 18.9 16.1 10.20 29.5 13.2 9.9
InstructNav [long2024instructnav] - - - - 7.44 31.5 25.2 19.1
NaVid [zhang2024navid] - - - - 6.74 36.3 26.6 20.8
ETPNav [an2024etpnav] 6.65 49.1 40.9 36.1 7.01 36.2 28.9 24.2
g3D-LF [wang2025g3d] - - - - 6.50 41.6 34.4 23.8

### v2

@@ IV-A2 Framework of the RSMP
a. Language Encoding Module: To capture the semantic structure of the instruction $I$, a pretrained language model such as BERT [15] is utilized. Let $\psi(\cdot)$ be the linear transformation layer, $\phi(\cdot)$ the embedding layer, and $\mathcal{F}_{I}(\cdot)$ the transformer encoder. The resulting language feature representation $Q_{s}\in\mathbb{R}^{L\times d_{h}}$ is computed as $Q_{s}=\mathcal{F}_{I}(\psi_{i}(I)+\phi_{i}(P_{I}))$, where $L$ denotes the number of the instruction tokens, $d_{h}$ is the hidden dimension of the feature embeddings, and $P_{I}$ denotes the absolute positional indices of the tokens. When the causal learning module is applied, the text content causal representations become $G_{s}=LN[\psi_{s}(BA(Q_{s},Z_{k}))]$, where $LN$ denotes the layernorm function, and $Z_{k}$ is the textual confounder features.
b. Panoramic Visual Encoding Module: To comprehensively perceive its surroundings, the agent processes a panoramic observation $E$ composed of 12 perspective sub-images. Each sub-image is first encoded by the pre-trained visual encoder $\mathcal{F}_{V}(\cdot)$ (like CLIP [50]). The relative orientation of each sub-image is represented by a directional embedding $\gamma_{im}={(\sin\theta_{i},\cos\theta_{i},\sin\beta_{i},\cos\beta_{i})}_{i=1}^{12}$. A multi-layer self-attention encoder ($\mathcal{F}_{\text{SA}}$) then fuses the multi-view features to enable cross-view interaction, obtaining $Q_{v}\in\mathbb{R}^{12\times d_{h}}$:
@@
c. Global-Local Cross-Modal Localization Module: In widely used VLN-CE hierarchical frameworks [3, 66, 67], cross-modal fusion is typically performed only at the global level between the instruction and the historical topology representation. However, this design may overlook fine-grained local context. Motivated by the effectiveness of dual local–global architectures in discrete settings [11, 69], we introduce a local cross-modal fusion branch to enhance detailed environmental grounding and action-level decision-making.
@@ IV-B1 Panoramic Visual Prompt (PVP)
To enable holistic spatial understanding, different from prior methods that rely on discrete sub-views or sequential video streams [9, 78, 13], we directly feed the full panoramic observation into the MLLM. Candidate waypoints are projected onto the panorama to visually encode their spatial positions, thereby enhancing the alignment between language instructions and navigational intent. Additionally, we observe that the lateral edges of the equirectangular panorama correspond to the agent’s backward view, which can mislead MLLMs into interpreting rear-view regions as lateral surroundings. Since navigation reasoning primarily depends on the forward field of view, we introduce a semi-transparent mask that attenuates the visual saliency of rear-view regions while preserving global scene continuity. This design reduces distractive cues from the backward view and helps the model concentrate on task-relevant front-view information. Our design is theoretically motivated by recent advances in attention prompting [74], which demonstrate that explicit visual prompts or region-wise masking can steer the attention distribution of large vision-language models and improve spatial reasoning. An illustration of PVP and the masking effect is shown in Fig. 3.
@@ IV-B3 Causal chain-of-thought
To enhance reasoning and interpretability, we apply chain-of-thought (CoT) prompting [70] to guide the RLMR to reason through five steps: trajectory status, action planning, visual grounding, suggestion evaluation, and final decision, with an assigned confidence score $\mathbf{p}^{\text{RLMR}}$. The complete prompt template is provided in the appendix.
@@ IV-C Uncertainty-Aware Collaboration Mechanism (UCM)
Existing works demonstrated that RSMP, operating within a panoramic-view hierarchical structure, consistently outperforms RLMR, even when the latter is fine-tuned on the in-domain VLN dataset [59, 79, 78]. While RLMR has the great potential to offer greater transparency and reflective reasoning, it is more prone to errors in long-horizon embodied reasoning tasks. To effectively integrate the strengths of both agents, we propose an Uncertainty-Aware Collaboration Mechanism (UCM), which leverages RSMP’s uncertainty estimates to adaptively mediate decision fusion with RLMR.
@@ IV-C1 Conformal Prediction for Uncertainty Quantification
To quantify RSMP’s prediction uncertainty, we adopt conformal prediction (CP) [52, 31], a distribution-free framework that provides statistically calibrated confidence measures with formal coverage guarantees. CP constructs a prediction set $\mathcal{C}(\mathcal{X})\subseteq\mathcal{Y}$ such that, in expectation over the data distribution, the true label $\hat{y}$ lies within $\mathcal{C}(\mathcal{X})$ with probability at least $1-\epsilon$. This coverage guarantee is statistically derived from the calibration set and is not enforced during evaluation or test inference. Specifically, $\mathcal{X}$ denotes the input observation space (comprising visual observations, language instructions, and navigation history), $\mathcal{C}(\cdot)$ denotes a set of possible output candidates, $\mathcal{Y}$ represents the action label space (possible waypoint candidates), and $\epsilon\in[0,1]$ is a user-specified miscoverage rate (i.e., the allowed probability of the prediction set not containing the true action).
@@ V-A1 Panoramic-RGBD end-to-end waypoint predictor
Most prior hierarchical VLN methods [37, 3] adopt an end-to-end waypoint predictor. The model takes RGB $\{v_{i}\}_{i=1}^{12}$ and depth $\{d_{i}\}_{i=1}^{12}$ panoramas, and encodes them using pre-trained visual backbones (denoted as $VE(\cdot)$ and $DE(\cdot)$, respectively). The visual embeddings are then processed by a Transformer encoder ($\mathcal{F}_{\text{SA}}$). A classifier ($\phi_{c}(\cdot)$) outputs a heatmap over a discretized 2D space defined by 120 angles (at $3\degree$ resolution) and 12 distances (from 0.25m to 3.0m in 0.25m increments). Finally, non-maximum suppression (NMS) is applied to the heatmap to extract the top-P candidate waypoints.
@@ V-B1 RGB-D-based point-Goal navigation network
For the local navigation controller, ETPNav [3] uses a “rotate-then-forward” strategy with a Tryout heuristic: if repeated FORWARD actions fail to change position, the agent rotates through preset angles $\Delta\Theta^{try}$ and retries. While effective in simulation, this trial-and-error approach is unsafe for real-world use. It often fails in scenarios with elevation changes, such as approaching a target atop stairs from the side. To address this, we adopt DDPPO [72], an end-to-end policy that maps RGB-D inputs and goals to actions, improving obstacle avoidance and deadlock handling in simulation.
@@ V-B2 SLAM-based point-goal navigation algorithm
LiDAR-based SLAM systems are widely used for robust real-world navigation, offering mapping, localization, and path planning in unknown environments without pre-built maps. Although DDPPO performs well in simulation, prior work [22] shows limited transferability to the real world. To ensure reliable physical deployment, we adopt an online SLAM-based pipeline using ROS move_base for waypoint navigation. This hybrid approach bridges the sim-to-real gap and ensures safe, efficient traversal.
@@ VI-A1 Datasets
The proposed method is evaluated on two widely used VLN-CE benchmarks: R2R-CE and REVERIE-CE, which are adapted from the discrete-path datasets [5, 47]. The R2R-CE dataset consists of real-world scanned environments across 90 buildings, with an average length of 32 words. It is split into train (10,819 pairs, 61 scenes), val seen (same buildings as training, 778 pairs and 61 scenes), val unseen (different buildings, 1,839 pairs and 11 scenes), and test unseen (3,408 pairs and 18 scenes). The test set is evaluated via an online server for fair benchmarking. The REVERIE-CE dataset includes 10,318 panoramas from 86 buildings and 21,702 crowd-sourced instructions, averaging 18 words per instruction.
TABLE I: Comparison with other SoTA methods on the R2R-CE dataset [38]. ‘-’ indicates the unavailable statistics. ‘Pano’ denotes whether panoramic visual input is used. $\dagger$ means the usage of the additional data beyond the VLN.
@@
I CM2 [21] 12.05 6.10 51 43 35 11.54 7.02 42 34 28 13.90 7.70 39 31 24
CWP-CMA [29] 11.47 5.20 61 51 45 10.90 6.20 52 41 36 11.85 6.30 49 38 33
CWP-RecBERT [29] 12.50 5.02 59 50 44 12.23 5.74 53 44 39 13.51 5.89 51 42 36
GridMM [68] 12.69 4.21 69 59 51 13.36 5.40 57 50 46 13.31 5.55 57 49 45
DreamWalker [56] 11.60 4.09 66 59 48 11.30 5.53 59 49 44 11.80 5.48 57 49 44
ScaleVLN [69] - - - - - - 4.80 - 55 51 - 5.11 - 55 50
ETPNav [3] 11.78 3.95 72 66 59 11.99 4.71 65 57 49 12.87 5.12 63 55 48
BEVBert [2] - - - - - - 4.57 67 59 50 - 4.70 67 59 50
HNR [67] 11.79 3.67 76 69 61 12.64 4.42 67 61 51 13.03 4.81 67 58 50
UnitedVLN [14] - 3.30 78 70 62 - 4.29 70 62 51 - 4.69 68 59 49
g3D-LF [66] - - - - - - 4.53 68 61 52 - 4.78 68 58 51
II InstructNav [45] - - - - - 6.74 6.04 42 30 22 - - - - -
NaVid [77] 10.27 5.47 52 43 38 7.63 5.47 49 37 35 - - - - -
Uni-NaVid [76] 10.04 4.45 65 58 53 9.71 5.58 53 47 43 - - - - -
NaVILA [13] - - - - - - 5.37 58 50 46 - - - - -
NaVILA [13] - 4.82 66 58 53 - 5.22 63 54 49 - - - - -
StreamVLN [71] - - - - - - 5.43 63 53 47 - - - - -
StreamVLN [71] - 4.53 68 62 58 - 4.98 64 57 52 - - - - -
@@ VI-A3 Implementation details
Our VLN-CE implementation builds on ETPNav [3]. The RSMP model is initialized with ScaleVLN [69] fine-tuned weights, with CLIP-H/14 [50] for image features and learnable backdoor and frontdoor causal modules from GOAT [63]. Depth information is incorporated via a pre-trained ResNet-50 [72]. Training in Habitat uses a batch size of 16 for 15K iterations with AdamW ($2\times 10^{-5}$), only on the standard R2R training split, and the best checkpoint is selected based on SR and SPL on validation-unseen. Compared with other LLM-based fine-tuning strategies that require large multi-GPU clusters, our method trains efficiently on only one single NVIDIA L40 GPU. REVERIE-CE is fine-tuned from the best R2R-CE checkpoint with identical settings. The RLMR deploys the open-source Qwen2.5-VL-72B [6] locally via VLLM [39] on 4 A800 GPUs.
@@ VI-B1 Results on R2R-CE
Table I compares CLASH with current SoTA methods on the R2R-CE dataset. For clarity, we categorize existing approaches into three types: Type-I includes conventional end-to-end task-specific models that are fully fine-tuned on the R2R-CE dataset. Type-II represents recent efforts to incorporate large models into VLN-CE. While these large-model approaches perform well in monocula settings, they still fall short of panoramic task-specific models, a trend also observed in discrete VLN benchmarks [79, 78]. Type-III refers to our proposed collaborative small-large framework, CLASH. It shows that CLASH outperforms existing methods across splits and metrics. Particularly, on the test unseen split, CLASH achieves the 1-st rank in the VLN-CE test leaderboard [^1] https://eval.ai/web/challenges/challenge-page/719/overview, surpassing g3D-LF [66] by a relative 13.79% in SR, and 13.73% in SPL, respectively. Compared to approaches using large vision-language models, CLASH demonstrates substantial gains in both seen and unseen environments. For instance, compared with NaVILA [13], CLASH increases relative SR by 25.86% on val-seen, and 20.37% on val-unseen, respectively.
@@ VI-B2 Results on REVERIE-CE
Table II compares our proposed CLASH method with other SoTA methods on the REVERIE-CE dataset. We mapped the REVERIE coordinates from the discrete environment into the Habitat simulator, enabling data collection, training, and model evaluation. It shows that CLASH outperforms previous methods on all metrics on both val-seen and val-unseen splits. For instance, it shows that on the val-unseen set, CLASH relatively outperforms g3d-LF [66] by 2.92% in SR and 11.76% in SPL, respectively. On the val-seen set, CLASH improves ETPNav [3] by relative 24.94% SR and 24.93% SPL, respectively. This has significantly demonstrated the effectiveness and robustness of our proposed method.
TABLE II: Comparison with other SoTA methods on the REVERIE-CE [47].
@@
Seq2Seq [38] 9.96 22.8 13.7 10.9 10.10 18.6 9.6 7.9
CMA-TF [42] 9.83 23.2 14.2 11.9 10.27 16.8 8.8 7.3
CMA-SF [42] 9.84 30.4 18.9 16.1 10.20 29.5 13.2 9.9
InstructNav [45] - - - - 7.44 31.5 25.2 19.1
NaVid [77] - - - - 6.74 36.3 26.6 20.8
ETPNav [3] 6.65 49.1 40.9 36.1 7.01 36.2 28.9 24.2
g3D-LF [66] - - - - 6.50 41.6 34.4 23.8